OpenStudy (anonymous):

what will the answer be to y/3 and it = 3

OpenStudy (aihberkhan):

Okay. Simply replace y with 3. When you do this you should get: \[3/3\] Now just solve! 3 divided by 3 is 1! So the answer is: \[1\]

OpenStudy (aihberkhan):

Hope this helped! Have a great day and Happy Holidays! :) Also a medal would be much appreciated! Just click best response next to my answer. Thank You! Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @yessicazepeda

OpenStudy (anonymous):

Do you mean y/3 where y=3 or y/3=3?

OpenStudy (anonymous):

the the 3rd option

OpenStudy (anonymous):

So\[\frac{ y }{ 3 }=3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what do you think is the first step to solving this one? Hint: The equation can also be written as \[(\frac{ 1 }{ 3 })y=3\]

OpenStudy (anonymous):

\[\frac{ y }{ 3 }=3\\ as\ this\ is\ \ a\ coditional\ equation , we\ need\ \to find\ variable\ y.\\ we\ know\ that\ multiply\ and\ division\ are\ inverse.\\ so\ , y= 3\times \ 3\\i.e.\ y=9\]

OpenStudy (anonymous):